XYZ, Inc. has budgeted sales for the first quarter of the next year to be 40,000 units. The inventory on hand at the beginning of quarter is 10,000 units. The desired ending inventory is 1000 units. Calculate the budgeted production for the first quarter.

  • Hospitality, Inc. has prepared the following direct materials purchases budget:

Month

Budgeted DM Purchases

June

$68,000

July

79,500

August

78,700

September

77,800

October

70,200

All purchases are paid for as follows: 10% in the month of purchase, 40% in the following month, and 50% two months after purchase. Calculate total budgeted cash payments made in October for purchases.
Answer #1

Answer:-

1) Budgeted sales - beginning inventory + desired ending inventory

= 40000-10000+1000

= 31000 units

2)payments in october

October (10%×70200) = 7020

September (40%×77800)=31120

August (50%×78700) = 39350

Total = 7020+31120+39350 = 77490